OpenStudy (anonymous):

How do I do multi-step operations with algrebraic fractions? I.e (1/x+1) + (2x/x^2-1) - (1/x-1)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge \frac{1}{x+1} + \frac{2x}{x^2-1} - \frac{1}{x-1}$$

OpenStudy (skullpatrol):

think about a common denominator...

OpenStudy (skullpatrol):

$$\Huge \frac{1}{x+1} + \frac{2x}{(x-1)(x+1)} - \frac{1}{x-1}$$

OpenStudy (anonymous):

Now when I have to use the LCD would it be (x-1)(x+1) ?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

Alright~ I'll try it on my own for now.

OpenStudy (anonymous):

I have done this so far: \[\frac{ x-1 }{ (x-1)(x+1)} + \frac{ 2x }{ (x-1)(x+1) } - \frac{ x+1 }{ (x-1)(x+1) }\]

OpenStudy (anonymous):

I heard that I need to do cross multiplication?

OpenStudy (skullpatrol):

That is sort of the idea...

OpenStudy (skullpatrol):

You need to use the rule for fractions that says you can multiply the numerator and denominator by the same number.

OpenStudy (skullpatrol):